Let M be the surface parametrized by T: (1, 0) R → R3 (u, v) = (ucov, usin 0,0 + 8}2+1] d) 1 Compute the mean curvature of
Answer #1

Ans: Given x lun) = (u cosu, usinu, V+ 2x du ( Cosu Sinu ax (-usin, u cosu, 1) SV E = Cos vt Sin²y + co 11 u-y + 4+1 E U4+1- ex le ( Sinu cosu ut u²1 ) - Cosv + Sinu uz (u) - textil Sinv -Cosu 2 [Sin u+1 U²1 + Tosv + Sinu u²+1 zlen - lexul Sint vtU21 X 름 2 (1²-1) (4²+1) Cosu M = 812 N = f Sinv) (sinv - Cosu v (Cosu + Sinu u ²-1 M = u²1 u 42+1 u 21 N= Yu . Ñ = (-ucosu) sU4+1 -2 (4²) - (U²+1)(4²1) 42 (4²4) M 9 (44+1) (4²+1) [ M = 4x +1 -24% -44 +2 +24% -2 2 [ 4²+1] [ +1-1] 2lu?+1)(9) su